# Quantum Langevin Dynamics for Optimization

### v1

@@ III.3 Convergence of QLD in the nonconvex case
$\braket{V(x)}_{t}-V(x^{*})\leq\frac{2}{r}kT+O(e^{-\eta rt}),$ (57)
@@
$\braket{V(x)}_{t}\leq\braket{H}_{t}\leq\frac{2}{r}kT+O(e^{-\eta rt}).$ (61)
@@
To achieve a desired accuracy $\varepsilon$, i.e., $\braket{V}_{t}-V(x^{*})\leq\varepsilon$ when $t\to+\infty$, we only need to set $kT=r\varepsilon/2$ and ensure that the approximation $\hbar\Omega\ll kT$ still holds.
IV Numerical results of time-independent QLD
In this section, we conduct numerical expertiments for time-independent QLD, i.e., all parameters remain constant throughout the evolution. In Section IV.1, we initially delve into the origin of dissipation, grounding our understanding in the intuitive concept of spontaneous emission. Section IV.1 also demonstrates that spontaneous emission can endow the system with the ability to transit to the ground state and achieve stability as the size of the heat bath increases. Section IV.2 primarily discusses the numerical methods employed to solving QLD. These technical details are essential for maintaining numerical stability during numerical computation. We then employ numerical experiments to validate the theoretical results concerning the quadratic potential from Section III.2, as detailed in Section IV.3. In the experiments of the quadratic potential, we found that the numerical results are in alignment with the theoretical predictions. This consistency validates the precision of our numerical experiments. Subsequently, roles of each parameter in QLD are individually discussed in Section IV.4. We discover that the parameter $\eta$ influences the speed of convergence in QLD, while the magnitudes of $\hbar$ and $T$ impact the system’s ability to escape from local minima. Finally, Section IV.5 presents a comparison of the performance between the time-independent QLD and its classical counterpart, the Fokker-Planck-Smoluchowski equation. This comparison demonstrates the consistency of the classical and quantum thermal effects.
@@ IV Numerical results of time-independent QLD
IV.1 Verification of the intuition: spontaneous emission
@@
We first verify that the system can still reach the ground state with high probability, when there is only a single two-energy-level oscillator and the Hamiltonian of heat bath is $H_{B}=\mathrm{diag}(1/2,3/2)\hbar\Omega_{1}$. The parameter setting and results are shown in Fig. 3, which implicates that there exists a frequency $\Omega_{1}=0.905\Omega_{0}$ such that the system can reach the ground state with high probability $|\langle\psi(t)|0\rangle|^{2}>0.9$.
Figure 3: Transition probability to the ground state, i.e., $|\langle\psi(t)|0\rangle|^{2}$. The quadratic case is a benchmark ($V(x)=\frac{1}{2}m\Omega_{0}^{2}x^{2},\Omega_{0}=w$), which exhibits characteristics of spontaneous emission (blue line). When $\Omega_{1}=0.905\Omega_{0}$, the system maximum probability of transition is larger than $0.9$ (orange line).
We increase the number of oscillators in heat bath, where $n=0,5,9,11,13$. The Hamiltonian of heat bath is $H_{B}=\sum_{n}\mathrm{diag}(1/2,3/2)\hbar\Omega_{n}$. For a physical system in the real world, the number of modes of frequency $\Omega_{n}$ should satisfy a specific distribution. For simplification, we choose the uniform distribution, namely the frequencies of oscillators being randomly chosen in the range $\Omega_{n}\in[0.7\Omega_{0},1.3\Omega_{0}]$. The results are shown in Fig. 4, which implicate that the probability of transition gradually stabilizes and stays in a high level with the increasing number of oscillators.
Figure 4: Transition probability when increasing the number of harmonic oscillators in heat bath. Since the frequencies of oscillators are randomly chosen, the results of this experiment are not unique. We choose the best one to verify our statements.
IV.2 Implementation of numerical experiments for QLD
@@
IV.3 Quadratic potentials
First, we verify whether the convergence values satisfy Eq. (53) and Eq. (54) by the following parameter setting  [^2] This parameter setting is also consistent with Refs. [75, 41, 52] after nondimensionalizing..
@@ IV.3 Quadratic potentials
Then, we try to verify whether the system reaches the ground state at the low temperature regime, stated at the end of Section III.2. The following parameter setting is used.
@@
Experiment Result: The evolutions are shown in Fig. 7 and Fig. 8. Still using same notations as previous, we have $\left\langle{V}\right\rangle_{f}=117.8211,~{}\left\langle{E_{k}}\right\rangle_{f}=117.8211$, $\left\langle{V}\right\rangle_{f}^{\prime}=117.4373,~{}\left\langle{E_{k}}\right\rangle_{f}^{\prime}=117.8249$, analytical solution of ground state potential energy and kinetic energy is $\frac{1}{4}\hbar\Omega=117.8211$. This consistency suggests that the system has reached the ground state. Denote $\rho_{f}$ to be the numerical solution of density matrix at the final moment $t=5$, and $\rho_{g}$ to be the analytical solution of the ground-state density matrix. We use fidelity and quantum relative entropy to evaluate the closeness between the numerical result and analytical ground state, which derives fidelity $F(\rho_{f},\rho_{g})=\left(\operatorname{Tr}\sqrt{\sqrt{\rho_{f}}\rho_{g}\sqrt{\rho_{f}}}\right)^{2}=0.9999902$ and quantum relative entropy (in Kullback-Leibler divergence) $D_{KL}(\rho_{f}\|\rho_{g})=\operatorname{Tr}(\rho_{f}(\ln(\rho_{f})-\ln(\rho_{g})))=6.3466\times 10^{-4}$, showing that the numerical result and analytical ground state are very close to each other.
@@
IV.4 Roles of each parameter
@@
IV.4.1 Roles of characteristic damping rate $\eta$
@@
Parameter Setting: Use different damping rates $\eta=0.5,~{}1,~{}2,~{}3,~{}5,~{}7,~{}10,~{}20,~{}30,~{}50,~{}100$, while other parameters are invariant. Resolution $n=128$, domain $x\in[-1,1]$, space step $\Delta x=\frac{2}{n}$, time step $\Delta t=0.05\Delta x^{2}$, total evolution time $t_{f}=10$, $\hbar=2.1108$, $kT=2072.6~{}(T=2500)$, $\Omega=300$, $m=1$, initial state $\psi_{0}(x)=\sqrt{p(x)}$.
@@
Figure 10: The time evolution of expectation values of potential. x-coordinate represents the time $t$, and y-coordinate represents the expectation value. From this figure, we can observe when $\eta$ is too small, the system has not reached the stationary distribution at $t=10$ still needing extra time to evolve. When $\eta$ is too large, the system reach the stationary distribution so fast that system cannot escape from local minimum completely and the convergence value of potential energy increases. Within a suitable range of $\eta\in[5,30]$, $\eta$ does not have a significant impact on the expectation value at $t=10$.
IV.4.2 Roles of $\hbar$ and $T$
@@
Parameter Setting: Change the reduced Planck constant $\hbar=1,~{}2.1108,~{}4,~{}8$,12,16,20,24,28,32,36,~{}12,~{}16,~{}20,~{}24,~{}28,~{}32,~{}36, 12 , 16 , 20 , 24 , 28 , 32 , 36, while other parameters are invariant in each case. Resolution $n=128$, domain $x\in[-1,1]$, space step $\Delta x=\frac{2}{n}$, time step $\Delta t=0.05\Delta x^{2}$, total evolution time $t_{f}=10$, $kT=0.0083~{}(T=0.01)$, $\eta=5$, $\Omega=300$, $m=1$, initial state: Gaussian distribution Eq. (65).
@@
Parameter Setting: Change the temperature $T=50,~{}100,~{}200,~{}500,~{}800,~{}1000,~{}2000,~{}2500,~{}3500,~{}5000$, $8000$ ($k=0.82904$), while other parameters are invariant. Resolution $n=128$, domain $x\in[-1,1]$, space step $\Delta x=\frac{2}{n}$, time step $\Delta t=0.05\Delta x^{2}$, total evolution time $t_{f}=10$, $\hbar=2.1108$, $\eta=20$, $\Omega=300$, $m=1$, initial distribution: Gaussian distribution Eq. (65).
@@
Parameter Setting: Conduct 100 cases, where $T=1000,1200,1400,1600,1800,2000,2200,2400,2600,2800$ and $\hbar=6,8,10,12,14,16,18,20,22,24$ respectively, while $\eta=20$ is fixed. We set an upper bound of the total evolution time $t_{f}\leq 100$, since it takes too long for some cases to reach the steady state. $\Omega=300$, $m=1$, initial state $\psi_{0}(x)=\sqrt{p(x)}$. Establish the criterion of steady state: let the $k$-th iteration of success probability be $\Pr^{(k)}$, defined by Eq. (64). Cease iteration when $\left|\Pr^{(k)}-\Pr^{(k-1)}\right|<10^{-7}$.
Experiment Result: As shown in Fig. 13, we observe that there exists a best temperature when $\hbar$ is fixed. Furthermore, although smaller $\hbar$ leads to higher success probability, the evolution time increases dramatically with respect to the decrease of $\hbar$. Due to the constraints of evolutionary economics, we cannot set $\hbar$ and $T$ to be too small.
Figure 13: We observe that there exists a best temperature when $\hbar$ is fixed. Furthermore, although smaller $\hbar$ leads to higher success probability, the evolution time increases dramatically with respect to the decrease of $\hbar$. Due to the constraints of evolutionary economics, we cannot set $\hbar$ and $T$ to be too small.
IV.5 Comparison with classical dynamics
@@
We compare QLD with FPE. In some landscapes, classical algorithms have demonstrated efficacy, and the advantages of quantum algorithms are not manifested. Hence, we choose a landscape that poses a comparatively challenging task for classical algorithms to serve as a basis for comparison, shown in Fig. 14. This function is called the Griewank function with form Eq. (105), which is a nonconvex function with many local minima approximating zero.
@@
V Time-dependent QLD
Based on the conclusions drawn in Section IV.5, QLD does not demonstrate a significant advantage in situations where classical algorithms already perform well. Conversely, in landscapes where classical algorithms exhibit shortcomings, QLD shows only marginal superiority. This implies that we have not fully leveraged the advantages inherent in quantum physics. Consequently, we propose a time-dependent QLD to address this limitation.
@@ V Time-dependent QLD
Section V.2 introduces process for selecting nonconvex objective functions. These functions are categorized into five groups, each with distinct features. In Section V.3, we provide an example of a specific time-dependent QLD for the nonconvex case by specifying the functions $\hbar(t)$ and $T(t)$.
We compare QLD with other state-of-the-art algorithms, including Quantum Hamiltonian Descent (QHD) [32], Quantum Adiabatic Algorithm (QAA), Stochastic Gradient Descent (SGD) and Nesterov’s accelerated gradient descent (NAGD) for different nonconvex objective functions, in Section V.4 and Section V.5. The first three algorithms are quantum algorithms, and the last two are classical algorithms. All quantum algorithms are simulated on classical computers to visualize the evolution of probability distributions. Finally, we summarize the results of comparison and put forward a three-phase diagram to explain the superiority of QLD in the nonconvex case, see Section V.6.
V.1 Convergence of time-dependent QLD
@@ V.1.1 Time-dependent $\eta(t)$
V.1.2 Time-dependent $\hbar(t)$
In the numerical experiments of time-independent $\hbar$, we observe that larger $\hbar$ can reinforce the effect of quantum tunneling to escape from local minimums, however, it also makes the system incapable of concentrating on the minimum. In order to solve this problem, we prove the following corollary of Theorem 1 with monotonically decreasing $\hbar(t)$ with respect to $t$.
@@ V.1.2 Time-dependent $\hbar(t)$
$\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}=$ $\frac{\mathrm{d}\ln\hbar}{\mathrm{d}t}\cdot\left\langle{p^{2}/m}\right\rangle_{t}+\left(\frac{\hbar^{2}\mu^{2}}{m}+\frac{\eta^{2}}{4\mu^{2}}\langle\nabla^{2}V\rangle_{t}\right)$
$-\eta\langle x\cdot\nabla V+p^{2}/m\rangle_{t}.$ (72)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}\leq$ $\frac{\mathrm{d}\ln\hbar}{\mathrm{d}t}\cdot\left\langle{p^{2}/m}\right\rangle_{t}$
$+\eta\cdot\left(2kT-\langle V+p^{2}/(2m)\rangle_{t}\right).$ (73)

### v2

@@ 3.4 Convergence of QLD in the quasar-convex case
$\left\langle{V(x)}\right\rangle_{t}-V(x^{*})\leq\frac{2}{r}kT+O(e^{-\eta rt}),$ (65)
@@
$\left\langle{V(x)}\right\rangle_{t}\leq\left\langle{H}\right\rangle_{t}\leq\frac{2}{r}kT+O(e^{-\eta rt}).$ (69)
@@
To achieve a desired accuracy $\varepsilon$, i.e., $\left\langle{V}\right\rangle_{t}-V(x^{*})\leq\varepsilon$ when $t\to+\infty$, we only need to set $kT=r\varepsilon/2$ and ensure that the approximation $\hbar\Omega\ll kT$ still holds.
4 Numerical results of time-independent QLD
In this section, we conduct numerical expertiments for time-independent QLD, i.e., all parameters remain constant throughout the evolution. In Section 4.1, we initially delve into the origin of dissipation, grounding our understanding in the intuitive concept of spontaneous emission. Section 4.1 also demonstrates that spontaneous emission can endow the system with the ability to transit to the ground state and achieve stability as the size of the heat bath increases. Section 4.2 primarily discusses the numerical methods employed to solving QLD. These technical details are essential for maintaining numerical stability during numerical computation. We then employ numerical experiments to validate the theoretical results concerning the quadratic potential from Section 3.3, as detailed in Section 4.3. In the experiments of the quadratic potential, we found that the numerical results are in alignment with the theoretical predictions. This consistency validates the precision of our numerical experiments. Subsequently, roles of each parameter in QLD are individually discussed in Section 4.4. We discover that the parameter $\eta$ influences the speed of convergence in QLD, while the magnitudes of $\hbar$ and $T$ impact the system’s ability to escape from local minima. Finally, Section 4.5 presents a comparison of the performance between the time-independent QLD and its classical counterpart, the Fokker-Planck-Smoluchowski equation. This comparison demonstrates the consistency of the classical and quantum thermal effects.
@@ 4 Numerical results of time-independent QLD
4.1 Verification of the intuition: spontaneous emission
@@
(a) A single two-energy-level oscillator in heat bath.
(b) Increasing the number of oscillators in heat bath.
Figure 3: Spontaneous emission. (a) Transition probability to the ground state, i.e., $|\langle\psi(t)|0\rangle|^{2}$. The quadratic case is a benchmark ($V(x)=\frac{1}{2}m\Omega_{0}^{2}x^{2},\Omega_{0}=w$), which exhibits characteristics of spontaneous emission (blue line). When $\Omega_{1}=0.905\Omega_{0}$, the system maximum probability of transition is larger than $0.9$ (orange line). (b) Transition probability when increasing the number of harmonic oscillators in heat bath. Since the frequencies of oscillators are randomly chosen, the results of this experiment are not unique. We choose the best one to verify our statements.
@@
We first verify that the system can still reach the ground state with high probability, when there is only a single two-energy-level oscillator and the Hamiltonian of heat bath is $H_{B}=\mathrm{diag}(1/2,3/2)\hbar\Omega_{1}$. The parameter setting and results are shown in Fig. 3a, which implicates that there exists a frequency $\Omega_{1}=0.905\Omega_{0}$ such that the system can reach the ground state with high probability $|\langle\psi(t)|0\rangle|^{2}>0.9$.
We increase the number of oscillators in heat bath, where $n=0,5,9,11,13$. The Hamiltonian of heat bath is $H_{B}=\sum_{n}\mathrm{diag}(1/2,3/2)\hbar\Omega_{n}$. For a physical system in the real world, the number of modes of frequency $\Omega_{n}$ should satisfy a specific distribution. For simplification, we choose the uniform distribution, namely the frequencies of oscillators being randomly chosen in the range $\Omega_{n}\in[0.7\Omega_{0},1.3\Omega_{0}]$. The results are shown in Fig. 3b, which implicate that the probability of transition gradually stabilizes and stays in a high level with the increasing number of oscillators.
4.2 Implementation of numerical experiments for QLD
@@
4.3 Quadratic potentials
First, we verify whether the convergence values satisfy Eq. (60) and Eq. (61) by the following parameter setting. [^1] This parameter setting is also consistent with Refs. [31, 29, 30] after nondimensionalizing.
@@ 4.3 Quadratic potentials
Then, we try to verify whether the system reaches the ground state at the low temperature regime, stated at the end of Section 3.3. The following parameter setting is used.
@@
Experiment Result: The evolutions are shown in Fig. 6 and Fig. 7. Still using same notations as previous, we have $\left\langle{V}\right\rangle_{f}=117.8211,~{}\left\langle{E_{k}}\right\rangle_{f}=117.8211$, $\left\langle{V}\right\rangle_{f}^{\prime}=117.4373,~{}\left\langle{E_{k}}\right\rangle_{f}^{\prime}=117.8249$, analytical solution of ground state potential energy and kinetic energy is $\frac{1}{4}\hbar\Omega=117.8211$. This consistency suggests that the system has reached the ground state. Denote $\rho_{f}$ to be the numerical solution of density matrix at the final moment $t=5$, and $\rho_{g}$ to be the analytical solution of the ground-state density matrix. We use fidelity and quantum relative entropy to evaluate the closeness between the numerical result and analytical ground state, which derives fidelity $F(\rho_{f},\rho_{g})=\left(\mathrm{Tr}\sqrt{\sqrt{\rho_{f}}\rho_{g}\sqrt{\rho_{f}}}\right)^{2}=0.9999902$ and quantum relative entropy (in Kullback-Leibler divergence) $D_{KL}(\rho_{f}\|\rho_{g})=\mathrm{Tr}(\rho_{f}(\ln(\rho_{f})-\ln(\rho_{g})))=6.3466\times 10^{-4}$, showing that the numerical result and analytical ground state are very close to each other.
@@
4.4 Roles of each parameter
@@
4.4.1 Roles of characteristic damping rate $\eta$
@@
Figure 9: The time evolution of expectation values of potential. x-coordinate represents the time $t$, and y-coordinate represents the expectation value. From this figure, we can observe when $\eta$ is too small, the system has not reached the stationary distribution at $t=10$ still needing extra time to evolve. When $\eta$ is too large, the system reaches the stationary distribution so fast that system cannot escape from local minimum completely and the convergence value of potential energy increases. Within a suitable range of $\eta\in[5,30]$, $\eta$ does not have a significant impact on the expectation value at $t=10$.
Parameter Setting: Use different damping rates $\eta=0.5,~{}1,~{}2,~{}3,~{}5,~{}7,~{}10,$ ${}20,~{}30,~{}50,~{}100$, while other parameters are invariant. Resolution $n=128$, domain $x\in[-1,1]$, space step $\Delta x=\frac{2}{n}$, time step $\Delta t=0.05\Delta x^{2}$, total evolution time $t_{f}=10$, $\hbar=2.1108$, $kT=2072.6~{}(T=2500)$, $\Omega=300$, $m=1$, initial state $\psi_{0}(x)=\sqrt{p(x)}$.
@@
4.4.2 Roles of $\hbar$ and $T$
@@
Parameter Setting: Change the reduced Planck constant $\hbar=1,~{}2.1108,~{}4,~{}8,~{}12,~{}16,~{}20,~{}24,~{}28,$ ${}32,~{}36$, while other parameters are invariant in each case. Resolution $n=128$, domain $x\in[-1,1]$, space step $\Delta x=\frac{2}{n}$, time step $\Delta t=0.05\Delta x^{2}$, total evolution time $t_{f}=10$, $kT=0.0083~{}(T=0.01)$, $\eta=5$, $\Omega=300$, $m=1$, initial state: Gaussian distribution Eq. (73).
@@
Parameter Setting: Change the temperature $T=50,~{}100,~{}200,~{}500,~{}800,~{}1000,~{}2000,~{}2500,~{}3500,$ ${}5000,~{}8000$ ($k=0.82904$), while other parameters are invariant. Resolution $n=128$, domain $x\in[-1,1]$, space step $\Delta x=\frac{2}{n}$, time step $\Delta t=0.05\Delta x^{2}$, total evolution time $t_{f}=10$, $\hbar=2.1108$, $\eta=20$, $\Omega=300$, $m=1$, initial distribution: Gaussian distribution Eq. (73).
@@
Parameter Setting: Conduct 100 cases, where $T=~{}1000,~{}1200,~{}1400,~{}1600,~{}1800,~{}2000,~{}2200,$ ${}2400,~{}2600,~{}2800$ and $\hbar=6,~{}8,~{}10,~{}12,~{}14,~{}16,~{}18,~{}20,~{}22,~{}24$ respectively, while $\eta=20$ is fixed. We set an upper bound of the total evolution time $t_{f}\leq 100$, since it takes too long for some cases to reach the steady state. $\Omega=300$, $m=1$, initial state $\psi_{0}(x)=\sqrt{p(x)}$. Establish the criterion of steady state: let the $k$-th iteration of success probability be $\Pr^{(k)}$, defined by Eq. (72). Cease iteration when $\left|\Pr^{(k)}-\Pr^{(k-1)}\right|<10^{-7}$.
Experiment Result: As shown in Fig. 12, we observe that there exists a best temperature when $\hbar$ is fixed. Furthermore, although smaller $\hbar$ leads to higher success probability, the evolution time increases dramatically with respect to the decrease of $\hbar$. Thus, it is necessary to balance the cost of evolution time and the success probability of the algorithm.
Figure 12: In the double well landscape, there exists a best temperature when $\hbar$ is fixed. Furthermore, although smaller $\hbar$ leads to higher success probability, the evolution time increases dramatically with respect to the decrease of $\hbar$.
4.5 Comparison with classical dynamics
@@
We compare QLD with FPE. In some landscapes, classical algorithms have demonstrated efficacy, and the advantages of quantum algorithms are not manifested. Hence, we choose a landscape that poses a comparatively challenging task for classical algorithms for comparison, shown in Fig. 13. This function is called the Griewank function with form Eq. (114), which is a nonconvex function with many local minima approximating zero.
@@
5 Time-dependent QLD
Based on the conclusions drawn in Section 4.5, QLD does not demonstrate a significant advantage in situations where classical algorithms already perform well. Conversely, in landscapes where classical algorithms exhibit shortcomings, QLD shows only marginal superiority. This implies that we have not fully leveraged the advantages inherent in quantum physics. Consequently, we propose a time-dependent QLD to address this limitation.
@@ 5 Time-dependent QLD
Section 5.2 introduces process for selecting nonconvex objective functions. These functions are categorized into five groups, each with distinct features. In Section 5.3, we provide an example of a specific time-dependent QLD for the nonconvex case by specifying the functions $\hbar(t)$ and $T(t)$.
We compare QLD with other state-of-the-art algorithms, including Quantum Hamiltonian Descent (QHD) [53], Quantum Adiabatic Algorithm (QAA), Stochastic Gradient Descent (SGD) and Nesterov’s accelerated gradient descent (NAGD) for different nonconvex objective functions, in Section 5.4 and Section 5.5. The first three algorithms are quantum algorithms, and the last two are classical algorithms. All quantum algorithms are simulated on classical computers to visualize the evolution of probability distributions. Finally, we summarize the results of comparison and put forward a three-phase diagram to explain the superiority of QLD in the nonconvex case, see Section 5.6.
5.1 Convergence of time-dependent QLD
@@ 5.1.1 Time-dependent $\eta(t)$
5.1.2 Time-dependent $\hbar(t)$
In the numerical experiments of time-independent $\hbar$, we observe that larger $\hbar$ can reinforce the effect of quantum tunneling to escape from local minima, however, it also makes the system incapable of concentrating on the minimum. In order to solve this problem, we prove the following corollary of Theorem 1 with monotonically decreasing $\hbar(t)$ with respect to $t$.
@@ 5.1.2 Time-dependent $\hbar(t)$
$\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}=$ $\frac{\mathrm{d}\ln\hbar}{\mathrm{d}t}\cdot\left\langle{p^{2}/m}\right\rangle_{t}+\left(\frac{\hbar^{2}\mu^{2}}{m}+\frac{\eta^{2}}{4\mu^{2}}\langle\nabla^{2}V\rangle_{t}\right)-\eta\langle x\cdot\nabla V+p^{2}/m\rangle_{t}.$ (80)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}\leq$ $\frac{\mathrm{d}\ln\hbar}{\mathrm{d}t}\cdot\left\langle{p^{2}/m}\right\rangle_{t}+\eta\cdot\left(2kT-\langle V+p^{2}/(2m)\rangle_{t}\right).$ (81)